OpenStudy (anonymous):

what is the sum of the first 100 even numbers

OpenStudy (anonymous):

dorema, the sum, 2 + 4 + 6 + ... + 98 + 100 is an arithmetic series (i.e. the difference between consecutive terms is constant (here, it's 2)). You will have 50 terms in this series, your first term will be 2 (sometimes 0, but that won't change the sum), and your last will be 100. The formula for such a sum is\[s_n=\frac{n}{2}(a+l)\]where n is the number of terms, a is the first and l is the last (there's also another useful version which you can look up). Your sum is then\[s_{50}=\frac{50}{2}(2+100)=25 \times 102 = 2550\]

OpenStudy (anonymous):

thanks that was a BIG HELP!

OpenStudy (anonymous):

become a fan, then :D